### v2

# LaserEscape: Detecting and Mitigating Optical Probing Attacks

Abstract.
The security of integrated circuits (ICs) can be broken by sophisticated physical attacks relying on failure analysis methods. Optical probing is one of the most prominent examples of such attacks, which can be accomplished in a matter of days, even with limited knowledge of the IC under attack. Unfortunately, few countermeasures are proposed in the literature, and none have been fabricated and tested in practice. These countermeasures usually require changing the standard cell libraries and, thus, are incompatible with digital and programmable platforms, such as field programmable gate arrays (FPGAs). In this work, we shift our attention from preventing the attack to detecting and responding to it. We introduce LaserEscape, the first fully digital and FPGA-compatible countermeasure to detect and mitigate optical probing attacks. LaserEscape incorporates digital delay-based sensors to reliably detect the physical alteration of the fabric caused by laser beam irradiations in real time. Furthermore, as a response to the attack, LaserEscape deploys real-time hiding approaches using randomized hardware reconfigurability. It realizes 1) moving target defense (MTD) to physically move the sensitive circuity under attack out of the probing field of focus to protect secret keys and 2) polymorphism to logically obfuscate the functionality of the targeted circuit to counter function extraction and reverse engineering attempts. We demonstrate the effectiveness and resiliency of our approach by performing optical probing attacks on protected and unprotected designs on a 28-nm FPGA. Our results show that optical probing attacks can be reliably detected and mitigated without interrupting the chip’s operation.
Deception, Laser Voltage Probe, Moving Target Defense, Optical Probing; Partial Reconfiguration; Side-Channel Attack;
journalyear: 2024copyright: acmlicensedconference: IEEE/ACM International Conference on Computer-Aided Design; October 27–31, 2024; New York, NY, USAbooktitle: IEEE/ACM International Conference on Computer-Aided Design (ICCAD ’24), October 27–31, 2024, New York, NY, USAdoi: 10.1145/3676536.3676822isbn: 979-8-4007-1077-3/24/10
1. Introduction
In various applications, the security of a computing machine relies on sensitive assets stored on Root-of-Trust integrated circuits (ICs). Examples of these on-chip assets include cryptographic keys, proprietary firmware, and intellectual property (IP). Attackers can extract these assets by gaining physical access to the devices where they are stored and mounting physical attacks. Attacks relying on failure analysis (FA) tools pose the biggest threat against ICs. Prominent examples of such attacks are the various forms of optical probing attacks (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021a, b) which enable the adversary to recover secret data and functions from chips. Illuminating a laser beam on transistors and measuring their states using the modulated backscattered laser beam is the main mechanism behind these optical probing attacks.
Deploying randomization in countermeasures, such as masking (Nikova et al., 2006; Groß et al., 2016), is a conventional method to mitigate conventional side-channel attacks, as it prevents the repetition and integration of the measurements over multiple clock cycles. However, randomness becomes ineffective if the adversary performs optical probing on halted circuits or data at rest (Krachenfels et al., 2021a, b). Moreover, many root/master keys, which cannot be masked, are vulnerable to optical probing during key caching from memory to registers (Lohrke et al., 2016; Krachenfels et al., 2021b).
A few prevention and detection-based countermeasures have been proposed in the literature to avert optical probing attacks. These countermeasures try to protect the design either directly from laser irradiation (Rahman et al., 2021; Parvin et al., 2023; Zhang et al., 2023) or detect tampering with the voltage/clock, which is required for certain optical probing attacks (Farheen et al., 2022). Such schemes deploy various logic styles and design techniques to either minimize the data/function-dependent leakage through the backscattered laser light (Rahman et al., 2021; Parvin et al., 2023) or to detect the generated heat illumination of the laser beam on logic cells (Zhang et al., 2023). Most of these solutions require the modification of standard cell libraries and are incompatible with legacy systems and field programmable gate arrays (FPGAs). Moreover, they are tailored toward a specific form of optical probing and, thus, cannot mitigate the threat in general. Finally, these schemes have yet to be fabricated or tested in practice.
The only FPGA-compatible sensing countermeasure that has been proposed so far deploys ring-oscillators (ROs) (Tajik et al., 2017a) to detect laser irradiation. However, such a solution has never been researched further due its high sensitivity to environmental conditions, which leads to high false alarms. In this work, we attempt to answer the following research questions: (1) Is it possible to confidently detect optical probing attempts in real-time using fully digital components of an FPGA? (2) What would be a suitable response upon detecting an optical probing attempt to deceive the adversary and not interrupt the IC’s operation?
Our Contribution. To answer the above questions, we present LaserEscape, which is capable of detecting and responding to optical probing attacks using the existing resources of mainstream FPGAs. To detect an optical probing attempt, we repurpose the recently proposed fully digital delay-based 1LUTSensor (Jayasinghe et al., 2023), which had been originally been designed for remote power analysis side-channel attacks. We found that the mechanism which allows 1LUTSensor to detect power fluctuations with only one look-up table (LUT) can also detect laser irradiation from the IC backside silicon. Afterward, to respond to the attack, we explore novel moving target defense and gate polymorphism strategies to deceive the adversary in real-time and render the attack ineffective. First, we demonstrate that how upon the detection of the attack, the partial reconfiguration (PR) feature of mainstream FPGAs and programmable SoCs can be utilized to dynamically randomize the placement of security-critical registers and moving them out from attacker’s probing field with a low overhead to protect secret keys. Second, we show how polymorphic logic gates can be realized on LUTs of FPGAs to zeroize the output of the logic function to prevent function extraction by the attacker. Finally, we apply LaserEscape to various circuits on a 28-nm FPGA and perform optical probing on them to assess the effectiveness of our solution.
Source Code Availability. We publicly publish our raw sensor data and make the source code of LaserEscape open source in https://github.com/vernamlab/LaserEscape
2. Technical Background
2.1. Optical Probing Attacks
Fig. 1 illustrates a high-level overview of optical probing attacks. As illustrated, a laser scanning microscope can be used to navigate through a target IC as well as capture dynamic transistor-level activities on it. Optical probing attacks often include extracting sensitive data from the target IC while it is operational (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021b, a). For this aim, a high-power incoherent light source (HIL) with a wavelength of 1.3 $\mu$m is utilized. In 1, the light beams are illuminated and passed through multiple optical mirrors, including scanning mirrors, before entering the objective lenses. The illumined beams are reflected from the IC, and are guided to the laser detector in 2 using splitters. By focusing the objective lens onto the region of interest (RoI), a 2D image of the target area can be captured and then be used for localizing physical components on the IC. Additionally by moving the stage different parts of the IC can be navigated.
Once focused, HIL beams can effectively pass through the silicon layer and, as shown in 3, reflect. The run-time alterations in the transistor-level components modulate reflected beams. In 4, these beams are captured, amplified, and prepared in the analog domain. As the received signals contain the activities of the scanned area, they can be used to localize components that operate at a certain target frequency using electro-optical frequency mapping (EOFM)( 5). Furthermore, by focusing on a particular node on the IC, high precision electro-optical probing (EOP)***When using a coherent light source, EOP is typically called Laser Voltage Probing (LVP), and EOFM is called Laser Voltage Imaging (LVI). can be mounted to estimate gate-level voltage fluctuations in pico-second time accuracy. At this point, the adversary can interpose the pattern image and activity signals to physically localize the targets on the IC as shown in 5. It is common that the adversary can perform multiple profiling attacks at this stage ( 6) by controlling I/O, global clock, and reset of the IC, and gain information about the target. For instance, by resetting the IC with a certain frequency or by injecting a particular pattern of inputs into the IC (Tajik et al., 2017b), the target elements could be identified using EOFM.
2.2. Sensor Implementations on FPGAs
In the case of laser attacks, such as optical probing, the colliding laser photons locally increase the temperature of the target wires and transistors on the chip. These local temperature variations can affect the propagation delays of electrical signals in delay-based sensors on FPGAs, such as ring-oscillators and time-to-digital converters (TDCs). ROs (Zhao and Suh, 2018) and TDCs (Schellenberg et al., 2021) have been extensively used for remote side-channel attacks on FPGAs, where they measure voltage fluctuations on the power delivery network (PDN) caused by adjacent intellectual property cores. While these sensors are often used for offensive side-channels, they can also be used for the detection of active attacks causing temperature and current variations. RO networks have been employed as the first attempt to detect optical probing attacks on FPGAs (Tajik et al., 2017a). The main drawbacks of using RO networks are their high power consumption and sensitivity to environmental conditions, leading to a high number of false alarms(Tajik et al., 2017a). Therefore, as described later, a more reliable TDC candidate will be chosen to detect optical probing attacks.
Figure 1. High-level overview of Optical Attacks and LaserEscape mitigation
2.3. Placement Randomization as MTD
To perform optical probing, the placement of the circuit under attack has to be constant for the duration of the attack, as the laser beam needs to stay for a minimum amount of time on each transistor to reveal the state of the transistor. Hence, if the placement of a circuit can be dynamically changed or randomized, optical probing becomes infeasible due to the reduced signal-to-noise ratio (SNR). To achieve randomized layouts for a circuit, the dynamic reconfiguration feature of mainstream FPGAs, such as partial reconfiguration (PR) or redundant logic and multiplexers on application specific ICs (ASICs) can be used. The PR feature of SRAM-based FPGAs represents a unique capability that allows for the dynamic alteration of specific sections of the FPGA while maintaining the uninterrupted operation of the remaining system. This functionality not only boosts the system’s flexibility but also optimizes power usage and enhances adaptability to evolving operational conditions or demands (Koch, 2012). Additionally, PR’s application in FPGAs has proven valuable in mission-critical settings where system downtime is impermissible and real-time processing is necessary (Koch et al., 2012).
AMD/Xilinx’s dynamic function exchange (DFX) introduces a technology for setting up PR areas within a static system, allowing users to allocate modules to these specific areas on FPGA fabrics (Xilinx, 2023c). However, the Vivado toolchain has several weaknesses, such as being too slow for real-time applications and the lack of support for bitstream relocation (Manev et al., 2022). On the other hand, the open-source tools, such as Byteman (Manev et al., 2022), improved the efficiency and speed of performing PR from an embedded processor, making it suitable for real-time applications, such as countermeasures against side-channel attacks during runtime.
Several countermeasures against side-channel attacks, like power and electromagnetic (EM) analysis, have already leveraged PR (Mentens et al., 2008; Güneysu and Moradi, 2011; Heyszl et al., 2012; Moradi and Mischke, 2013; Hettwer et al., 2019; Bow et al., 2020; Khan et al., 2021). Some of these approaches use PR to introduce timing jitter, which can help distort power traces and defeat power side-channel attacks. Other techniques involve dynamically relocating functional units to different regions of the FPGA through PR, dissociating the spatial correlation between EM emanations and the target logic. The capability for real-time reconfiguration of FPGAs through PR has also been explored to create side-channel resistant implementations for cryptographic primitives, such as (Monfared et al., 2024).
2.4. Gate-Level Polymorphism as MTD
Another approach to respond to an optical probing attempt, in the case of function extraction, is to transform the gate functionality. Such real-time function transformation can be obtained using polymorphic circuits. Polymorphic circuits are digital circuits with the ability to perform two or more logical operations, where the selected operation depends on external factors such as temperature, light, or supply voltage. Polymorphic circuits were first introduced in CMOS circuits (Stoica et al., 2001) using evolutionary algorithms. The classic example of a polymorphic circuit is a gate whose output functionality is OR when operating at high supply voltage, but switches to AND at low supply voltage. Further works have explored static design methodologies for polymorphic gates without the need for evolutionary algorithms (Bernard et al., 2021). These circuits can be used in CMOS circuits to change pipeline functionality with low area impact and are compatible with a wide variety of logic function combinations.
While most of the polymorphic gate realizations are carried out by custom cell libraries, multiplexers and switch boxes have also been used to design polymorphic elements (Falkinburg, 2011). In this work, we leverage the flexibility of LUTs within FPGAs for the implementation of polymorphic circuits. LUTs can be programmed with the union of the truth tables of two different logic functions, with one input to the LUT selecting which logical function is propagated to the output. By combining this LUT-based approach with external sensors, gate-level polymorphism can be achieved without the need for reprogramming LUTs or rerouting the circuit.
3. Threat Model and Attack Scenarios
For optical probing attacks, we presume that an adversary is equipped with FA tools and aims to extract a secret key or reverse engineer specific elements. The primary motivation for these attacks comes from the scenario where a single key is employed across all devices, such as in instances where firmware, bitstream, or logic encryption secures the proprietary design of a system. The vendor programs this key into the product before distribution to customers. Thus, acquiring the key from one device could compromise the security of the entire device family. Moreover, even if individual devices use different keys, the uniform layout across a device family means that once an adversary masters key extraction from one device, she can apply this knowledge to breach other devices within the same family. At the right side of the Fig. 1, attack scenarios are highlighted in 7. For this article we consider optical-based 1) key recovery attacks(Krachenfels et al., 2021b, a), and 2) function recovery attempts(Mehta et al., 2024).
To evaluate each of these scenarios, we consider suitable optical probing techniques. Specifically, we consider both EOP and EOFM for key recovery, and for function reversing, we consider an EOFM attack. Then, we deploy LaserEscape prototype on the target device and verify our mitigation methodology accordingly. As displayed in Fig. 1, our proposed detection and mitigation methodologies are highlighted with 8 and 9. We implement a low-overhead delay-based sensor in proximity to the target elements. We assume that the sensor is supplied by an internal clock source so that the adversary cannot tamper with it. Upon the illumination of the HIL at the protected physical zone, the physical characteristics of the target are temporary altered and sensed by the proposed detector in 8. This triggers the defense response on the DUT to modify the target element. In 9 the indicated elements are modified, removed, or redeployed based on the strategy of the defender in real-time.
4. Attack Detection
4.1. Sensor Requirements
As mentioned in Sect. 2.2 optical probing attacks induce temperature fluctuations within the chip. These localized variations alter the signal propagation delays of the transistors. An ideal sensor must satisfy multiple conditions to detect such attacks:
Large Spatial Coverage. Performing optical probing over a specific area only increases the temperature locally in the wires and transistors of that region. Therefore, an ideal sensor should provide coverage across the entire security-sensitive areas to detect localized temperature variations.
Large Temporal Coverage. The irradiation of individual gates or registers during an optical probing attempt is potentially an extremely rapid process (e.g., less than 1 ms). Hence, the sensor must have a higher time resolution and continuity to capture an attack successfully.
Reliability. The sensor must be immune to voltage and temperature variations, so to detect the attack under various environmental conditions reliably. Otherwise, the sensor could detect noise as an attack attempt and generate false alarms.
4.2. Sensor Architecture
Based on the above requirements, we will present the architecture of our delay-based sensors in this section. Our sensor design is inspired by the 1LUTSensor originally implemented for on-chip power analysis attacks implemented in (Jayasinghe et al., 2023), see Fig. 2. It consists of a clock signal connected to the data and clock inputs of a register simultaneously. Moreover, the delay of the clock and data paths can be tuned using PVT (Process, Voltage, Temperature) invariant delay elements, namely one or multiple IDELAYE2 elements found in the I/O blocks of the AMD 7 series FPGAs (or similar tapped delay elements in other FPGA families). The signal on the register’s data input path passes through a LUT, in which the path from the LUT inputs to the output can be configured for further adjustments.
At runtime, we set the delays of each path and select the input of the LUT to pass through such that the data pulse arrives a short moment after the setup time boundary, as shown in Fig. 3. In this case, in principle, the sensor output stays at 1. By increasing the delay of the LUT, the output moves toward metastablity. We can validate this behavior by iterative sampling the output of the register and observing more 0 occurrences. As a result, by finding the right tuning levels, the sensor could be stabilized by outputting 1 and getting triggered (outputting 0) when its delay is increased due to an increase in temperature.
The primary reason behind choosing this sensor is its size compared to other TDCs, in that it only requires one LUT and can be placed in any slice, satisfying the spatial coverage requirements. Furthermore, the utilization of these PVT-invariant elements for delay tuning makes the sensor immune to environmental conditions, satisfying its reliability. Finally, supplying the sensor a few tens of megahertz clock signal provides sufficient timing resolution less than the laser dwell time on a pixel.
A key difference in our implementation is using a much lower clock frequency. The original 1LUTSensor as described in (Jayasinghe et al., 2023) was developed to measure voltage fluctuations in a circuit running on the chip to perform power analysis attacks. As such, it ran as fast as it could on the FPGA fabric - 600MHz - to measure an AES circuit running at frequencies ranging from 12-120MHz with enough resolution to detect the effects of switching gates. We do not want such a precise resolution, since this would cause the circuit we wish to protect to add noise to our readings. In addition, the laser pulses for LVI are on the order of milliseconds, which means the fluctuations in the sensor may be of that size.
Figure 2. Block diagram of 1LUTSensor
4.3. Adjustable Delay Elements
Tapped delay elements are extensively utilized in FPGAs to facilitate signal delays. These elements are capable of adjusting the delay of a wire in run-time. Different delay blocks are available for different FPGA families. For instance, AMD provides IODELAY1, IODELAY2, and IODELAY3 blocks in Spartan 6, 7 Series and UltraScale/+ families, respectively, featuring different adjustable delay blocks. (Jayasinghe et al., 2023) Here, we describe the details of our sensor based on IDELAYE2 elements. However, the technique presented here can be extended to other delay blocks as well.
IDELAYE2 elements are 31-tap, wraparound, delay elements with a calibrated tap resolution (Xilinx, 2023). Each IDELAYE2 acts as a pass-through for a 1-bit signal with a 5-bit input for the delay configuration, where the higher value leads to a larger delay. Since the delay of the clock path is larger than that of the data path, a single IDELAYE2 might not provide sufficient delay. Therefore, a chain of IDELAYE2s for the clock path can be used as shown in Fig. 4.
(a) Constant 1 case
(b) Constant 0 case
(c) Metastable case
Figure 3. Timing diagrams of 1LUTSensor
In our design, the chain consists of $2^{n}$ IDELAYE2s, where $n$ is any positive integer, and the delay input is $n+5$ bits wide. The five least significant bits set the delay of one of the IDELAYE2s, and the $n$ most significant bits set how many of the rest have the maximum delay (11111). The rest have the minimum delay (00000). For example, in the 4-long chain shown in Fig. 4, $n=2$, so the tuning input is 7 bits wide. Similarly, if the delay were 1010111, one IDELAYE2 would have a delay of 10111, two IDELAYs would have the maximum delay, and one would have the minimum delay. Moreover, we can verify the behavior of these delay blocks by including them in series with a RO on an AMD Kintex 7 FPGA. In this case, we can keep track of the RO’s frequency based on the IDELAYE2 delay(s). Furthermore, by plotting the corresponding chain period as a function of delay values, we can determine the range and resolution of these delay blocks for fine-tuning the sensibility. Fig. 5 plots the period of an exemplary 11-stage RO chain with 8 IDELAYE2 blocks, showing $ps$ resolution for the sensor.
4.4. Sensor Tuning
In our sensor implementation, we use a standalone IDELAYE2 for the data path delay and an $m$-long IDELAYE2 chain for the clock path. Moreover, we use a select input to determine which input of the LUT goes through to the register. Hence, we establish a tune value, consisting of three values: 1) data-path delay, 2) clock-path delay, and 3) LUT select. For many combinations of data path delay and LUT select, there is a corresponding value for the clock path delay where the output is metastable. However, some combinations perform better since their accumulative zero counts are significantly smaller than others, which means they are metastable in a way that is closer to the constant zero case shown in Fig. 3(a). For every value of the data path delay, the binary search of clock delay for metastability is carried out with LUT select fixed at zero. If metastability is found, we search for adjacent clock delay values with every LUT select value for other metastable tunes. During this process, we keep track of the best tune, determined by the lowest maximum zero count over the $t_{sense}$ interval. At the end of the procedure, we extract the best tune.
Figure 4. Block diagram of a 4-long IDELAYE2 chain
Figure 5. Graphs of ring oscillator period vs delay in IDELAYE2
4.5. Sensor Output Post Processing
When the sensor’s output is metastable, yielding an output of $1$ with occasional zero pulses, the length and frequency of these pulses increase during optical probing due to the impact of increased temperature on delay lines. Hence, we utilize measurement circuitry to characterize this. First, we implement a zero counter, which counts the number of clock cycles for which the sensor value is zero over $t_{detect}$. Secondly, we employ a pulse counter, which counts the length of the zero pulses. Based on the readings of the counters, we connect the sensor output to a latch, which we use to trigger our chosen defenses: 1) partial reconfiguration or 2) polymorphic gates, which will be described in Sect 5.1 and Sect 5.2, respectively.
5. Attack Response
Real-time response against physical attacks can be addressed with different approaches in accordance to the threat model. For instance one can deploy a self-destructive mechanism to zeroize the sensitive data upon sensing the physical attacks (Cannon et al., 2023; Tada et al., 2021). In contrast to solutions which necessitates cell library modification at pre-silicon (Shen et al., 2018), we propose fully logic-level mitigation methodologies that provides non-interrupting circumvention in addition to existing CMOS friendly zeroization (Tada et al., 2021; Cannon et al., 2023). Specifically, in our method target circuitry can benefit from real-time reconfiguration to bypass the optical attack and almost uninterrupted, operate as normal. This is useful for deceiving the adversary (Tajik et al., 2017a). Furthermore, the full zeroization of the device including the entire memory contents incurs a huge area and latency for large designs, (Cannon et al., 2023; Srivastava and Ghosh, 2019) which can be replaced by partial zeroization. In addition to MTD, we incorporate a gate polymorphic design, particularly to defeat optical reverse engineering attempts. We design the polymorphic gates by inspiring the state-of-the-art methodology in (Cannon et al., 2023) to create a supply voltage-controlled gates and obfuscate the functionality of the target elements in real-time.
5.1. Moving Target Defense via PR
Figure 6. The Operation flow of the proposed Moving Target Defense in an FPGA.
Fig. 6 illustrates the design of our proposed MTD-based response for an optical attack on an FPGA. We consider an original hardware design which includes some sensitive data requiring protection. For simplicity, we assume the key registers $R_{1}$, $R_{2}$, and $R_{3}$ as the secret elements in our design. Using TCL, the targeted modules are specified by the user through predefined annotations in register-transfer level (RTL). At this stage, the static design in the original bitstream as well as sample partial files are generated. The constraints such as the allowable range for FPGA slices, regional locations, and possible range of flip-flops (FFs) (Xilinx, 2023b) are established for generating reconfiguration. This information, together with the generated bitstreams, are moved to the online phase of the framework. In this phase, a lightweight processing unit (PS) is employed within the SoC (or in an external processor) to execute reconfiguration. A secure, one-time pseudo random number generator (PRNG) (Tsoi et al., 2003; Monfared et al., 2020) is implemented to provide randomness to the PR-generator unit each time. The PR-generator includes a bitstream manipulator (Byteman(Manev et al., 2022) in our instance), uses the predefined constraint limits to randomized location (LOC) and shuffle the chosen registers. These two randomization introduce coarse and fine-grain randomness to the reconfiguration, respectively. Consequently, the corresponding partial reconfiguration bitstream is created based on the original bitstream. The FPGA is then reprogrammed as soon as the trigger signal from the sensor is received.
Inter-Register Hiding. The core functionality of the proposed online phase is bitstream manipulation aimed at generating one-time PR. This approach introduces real-time randomization, deviating from previous methods where bitstreams were stored in memory (Khan et al., 2021). This technique not only enhances the security level of the countermeasure but also reduces the memory usage of the PR files to a single bitstream. As depicted in Fig. 6, the PRNG unit on the processor generates one-time randomness, and the bitstream manipulator program (e.g., Byteman) collects this randomized constraint information to produce one-time bitstream. In the coarse grain randomization, one or multiple slices in the FPGA are considers. All the corresponding elements in the target slices (e.g., target FFs and LUTs) are identified in the bitstream. The bitstream manipulator activates the corresponding routing and elements on another slices in a randomized possible reconfigurable module on the fly and disables the original slices of the target module and the FPGA proceeds to reconfiguration. The reconfiguration process delay overhead is considerably less than EOFM/EOP integration time, leading to successful mitigation and deceiving the attacker. These timing are investigated in detail in Sect. 7.
Intra-Register Hiding. For fine-grain manipulation, we utilize DFX to implement PR as a hardware-based scrambling method to counter Laser Voltage Probing attacks. For this purpose, a high-level script (such as Python) is used to randomly permute target registers (e.g., the 3 bits of a master key shown in Fig. 6). The potential search space for such permutations is super-exponential ($n!$), which effectively defeat probing attacks. This sequence is then transferred to the bitstream manipulator program to be used as constraints within the bitstream codes. Unlike the previous method, this approach minimizes resource utilization by requiring only a single reconfigurable module for implementation. This efficiency arises because the only circuitry modification in this method involves the hardware referencing of the target flip-flops, which effectively alters the local internal routing in the target slice.
(a) As one 4-input LUT
(b) As two 3-input LUTs
Figure 7. Block diagram of a 4-input LUT implementing a 4-input AND gate
5.2. Gate Polymorphism
To implement a polymorphic gate, we exploit the LUTs on the FPGA. An $n$-input LUT consists of $2^{n}$ SRAM cells feeding into a $2^{n}$-to-1 multiplexer. One could formulate the truth table of an $n$-input logic function to the SRAM cells and use the select inputs as the inputs of the logic function, choosing which SRAM cell to pass on to the output. An example of a 4-input LUT implementing an AND gate is in Fig. 7(a). Alternatively, this 4-input LUT can be seen as two 3-input LUTs with a 2-to-1 multiplexer to select which input is passed through to the output, as shown in Fig. 7(b).
In general, for $n$-input LUTs, we can implement polymorphic $n-1$-input LUTs using a control input to change the logic function. For instance, we can exploit input $i3$ in Fig. 7(b) to switch the output between 1) The result of a 3-input AND and 2) constant zero. This idea can be generalized to implement other logic gates (i.e., XOR) easily by selecting proposer LUT inputs. Furthermore, by connecting the control input pin to the latch output of the sensor described in Sect. 4.2, we can modify any protected logic upon detecting optical probing attacks.
6. Experimental Setup
Figure 8. Measurement setup. (a) Objective lenses focusing on the target IC (b) High-level setup diagram.
Electrical and Optical Equipment. A Hamamatsu PHEMOS-X FA microscope was used to perform optical probing, see Fig. 8. It is equipped with a 1.3 $\mu$m high-power laser source used for our analysis. Additionally objective lenses of 5X/0.14NA, 20X/0.6NA, 50X/0.76NA, and 71X are employed in our experiments. Furthermore, optical zooms of 2x, 4x and 8x were higher magnifications. As illustrated in Fig. 8(a), in the EOFM process, the DUT is scanned by a laser through the use of galvanometric mirrors.
(a) timeline illustration of EOFM scan to extract targeting register values
(b) timeline illustration of EOFM scan to extract targeting register values while LaserEscape activated.
Figure 9. EOFM probing attack timeline on images captured during laser scanning from back-side of the FPGA.
Device Under Test. For our experiments we used Digilent’s Genesys 2 development board, equipped with a Kintex-7 FPGA from Xilinx/AMD (part number XC7K325T-2FFG900C). We developed a custom-made 3D printed stage to hold the target device and the controller system.
Communication. For the run-time configuration/communication of the target FPGA, we used a 3.3V modified Arduino Uno R3 communicating over I2C interface with the FPGA target. A compact circuitry on the FPGA part is deployed to handle I2C communication and commands from the user. We choose an I2C bus since the addressable communicated data makes adding post-processing circuits easy for different experiments.
Bitstream Reconfiguration. For the MTD implantation, once the target is initiated with the mitigation mechanism, we utilize a Raspberry Pi 3 Model B equipped with a Quad Core 1.2GHz Broadcom 64bit CPU for both partial reconfiguration and programming.
Analyzer and Controller Configuration. Illustrated in Fig. 8(a), we use a dedicated controller system, to send commands, program key registers, adjusting target clock, and handle I/O for experiments via I2C. The measurement system is a dedicated system synced with the PHEMOS-X to control the stage, laser configurations and display the digitized data captured from the detection RAC. We use Semishop Failure Analysis Software 4.0 to control the stage and capture the images and signals.
Optical Probing Configuration. For the EOFM and EOP analysis we used 20x objective lens to focus on the target circuitry. The EOFM target frequency in our experiments is set to 1.25MHz, meaning that we use the controlling system to alternate the global Rst of the target every $200\mu s$. The integration time for the EOFM scan speed is set to 1 ms/px with 1KHz of bandwidth. For the EOP, the probing resolution is set to 100ps with 10,000 iterations. With the aforementioned configurations, a complete EOFM scan in 1x zoom takes approximately 194 seconds and as for a complete integrated EOP it takes 24 seconds.
Sensor Configuration. After implementing the target RTL, we added the sensor in a neighbour SLICE of the target. In our sensor implementation, we used a chain of 8-long IDELAYE2 ($m=8$) for clock delay path. The tuning time interval is set to $t_{sense}=100ms$ and the zero counter interval window in test time is set to $t_{detect}=255$ clock cycles. We chose $m$ and $t_{detect}$ because they were the largest values we could read or write with a single byte over I2C bus. We chose $t_{sense}$ because to achieve a balance of sensitivity versus speed when conducting experiments.
7. Results
7.1. EOFM for Localizing Target Registers
For the EOFM attacks on the key registers, we follow the similar attack flow described in (Lohrke et al., 2016). Specifically, we mount a profiling attack on key registers implemented by flip-flop with clock enable and synchronous reset (FDRE(Xilinx, 2023a). We carry out a localization attack accurately by exploiting the reset signal of the DUT (Lohrke et al., 2016). In this attack, the adversary can localize the coarse-grain circuitry of interest by feeding the known values into the chip via I/O with a particular frequency. Then, by sweeping in a particular frequency bandwidth with the center frequency of I/O activities, she can effectively discover the target registers and the corresponding connecting circuitry near them on the chip. Fig. 9(a) shows the timeline of an EOFM scan on a target area. As shown, the target registers are accurately localized on the selected frequency (i.e., reset frequency) once the EOFM scanner line reaches the area.
Images captured from back-side FPGA in Fig. 9(a) show how the scanning is carried out (from left to right). As the horizontal scan line (across Y axis) moves, details of the internals of the FPGA are shown. The bright spots represent the electronic elements (e.g., registers) that actively operate on particular frequencies within the range of the EOFM scan. The target register areas, highlighted in a red box, are programmed registers that are localized by profiling. Similar to methods in (Lohrke et al., 2016), the profiling process consists of 1) programming the elements to fluctuate in a certain frequency, 2) activate/deactivate them and 3) analyze the optical activity extracted from EOFM. In our case shown in Fig. 9(a), 1) the target FDRE registers are programmed to reset with EOFM scanning frequency, 2) are set and clear in multiple iterations and 3) are inspected visually after EOFM scan in terms of the optical activity. The bright area within the red target area indicates that 1) the registers are fluctuating on the particular frequency and 2) they are set to one.
7.2. EOP for Register Values
Upon successful localization of the registers, we can then mount an EOP attack by focusing the HIL beam on each target FF. This process can be executed during the secret operation of the IC. For instance, if the state registers of a block cipher (e.g., AES) are probed, the keys are revealed easily. Fig 10 showcases a successful EOP attack on an implementation of shift-register. In this particular scenario, we focus the lens onto the registers (which are already localized via EOFM) and carry out the shift operation that is programmed on the FPGA. The probing process is synchronized with the operations on the FPGA via triggers. The probing experiment is performed during a particular interval with picosecond time accuracy to capture fast logical operations which results in voltage alterations in nodes on the chip. After multiple iterations and measurement averaging, the curves illustrated in Fig 10 are produced. Sometimes signal processing techniques, such as frequency band filtering and de-noising, are required to extract the clean states of the targets. This is due to the errors in placing and focusing the lens at the target element. As depicted in Fig 10, the shift register operations in REG 1 and REG 2 can be followed synchronously with the CLK-BUF rising edges.
Figure 10. EOFM and corresponding EOP waveforms on target registers.
7.3. EOFM for Function Recovery
As an extension to the key recovery attempt, we also demonstrate a function recovery attack via EOFM. Similar to previous probings, the adversary localizes the region of interest by exploiting the control signals in the frequency domain. By conducting a careful profiling procedure, a reverse-engineering approach is possible to discover the functionality of the building blocks. Fig. 11 serves as an example for a series of profiling EOFM patterns of registers in a function. It can be inferred that the correlation of the FFs patterns can be described by an XOR function among the 8 input FFs at the left side of the image and the output FFs at the right of the image. As indicated, the programmed functionality for this function is a simple bit-wise XOR operation where $c_{i}=a_{i}\oplus b_{i}$. By analysing the EOFM optical response from different input patterns, we can visually confirm the activation patterns for each individual register bits. By a simple observation at the first two rows of Fig. 11, one can suggest that the response of the neighbouring underlying registers are separated two by two. Particularly as highlighted by red squares, the optical patterns of $a_{3}a_{2}$ and $a_{1}a_{0}$ are physically separated. This might suggest the physical realizations of FFs in each slice on this particular FPGA architecture. Furthermore by analyzing the last row in Fig. 11, we can recover the functionality of the operations programmed for this target. Specifically, for $a_{1}a_{0}=0b01$, we can observe that that brightness pattern differs from $110b$ 11. With the same observation on $c_{1}c_{0}=0b10$ we can recover the functionality of the underlying operation in this analysis.
Figure 11. Set of EOFM images for function(XOR) recovery
7.4. Evaluating Sensor Trigger
In order to ensure the sensor operates properly, we empirically investigate the output of the sensor. As an example, Fig 12 displays the normalized zero counter output of the sensor in two test cases during a time window of $t_{detect}=255$ clock cycles. As it can be observed, while the laser is activated, the average of the zero counter samples is significantly larger compared to the scenario where the laser is idle. Furthermore, Fig 12 shows that in this case, the trigger threshold can be selected by a safe margin.
Figure 12. Sensor output against optical probing attempt.
Note that compared to the EOFM/EOP integration time, the highlighted time window where the sensor samples are collected is very short. Specifically, for a $f=100MHz$ operation frequency, the samples are collected during $256\times 10^{-8}=2.56\mu s$, which is considerably faster compared to the dwell time of the laser during probing.
7.5. Defeating Function Discovery Attacks
For this experiment, we deploy LaserEscape sensor on the target FPGA. We mount the same optical probing attack described in Sect. 7.3. In this experiment instead of a simple XOR gates for the target function we deployed the proposed polymorphic XOR gates in the target design. For the EOFM, Fig. 13 displays the input and output registers of the target function. As shown in the upper part of Fig. 13, the unprotected target reveals the XOR function for $in_{1}=0b0101$ and $in_{2}=0b1010$ which results in $out=0b1111$ highlighted on the right side of the figure. When the sensor detects laser emission on the FPGA, the polymorphic gate activates and clears the output of the XOR gates. This is highlighted in the lower part of Fig. 13.
Figure 13. EOFM images for function(XOR) recovery and polymorphic gate mitigation.
7.6. Defeating Register Recovery
To evaluate our method we organize the similar EOFM attack described in Sect. 7.1 in presence of LaserEscape. We target a set of 8 FDRE registers which are considered to be protected in the design. One can assume these registers hold the master key of a cryptographic algorithm (e.g., AES). Fig. 9(b) illustrated the timeline of the EOFM scan targeting the FDRE registers. Compared to the similar attack depicted in Fig.9(a), as the scanning laser line approaches the physical locations of the target FDREs, the sensor in LaserEscape is triggered and consequently the MTD via PR is carried out in real-time, leaving no trace of the FDRE targets once the laser arrives at the target destination. Note that the sensor circuitry is placed on a neighbouring slice to the target shown in Fig. 9(b) and is not illustrated in these captures.
8. Discussions
8.1. Overhead Analysis
In terms of the overhead introduced by LaserEscape, the 1LUTsensor is very efficient in terms of area overhead by employing a single LUT and utilizing 9 IDELAYE2 in the design, which is negligible. For the MTD, if 32-bit state registers of an AES block cipher is considered as the target, a total area overhead of 3.5% (compared to original AES) is estimated by Vivado toolkit in the case of choosing 8 Reconfigurable LOCs. The delay overhead caused by the PR process for the same implementation is measured to be $223\mu s$, which maintains a large safe gap against the laser dwell time per pixel for EOP and EOFM attacks.
8.2. Stability
As long as the ambient conditions and FPGA load remain the same, our sensor remains stable. We let the sensor run idle for about 21 hours, logging the sensor’s readings over zero counts. Our experiments, shown in Fig. 14, show that the maximum zero count had plateaued after less than 20 minutes, but the rolling average stayed at around zero for the rest of 21.5 hours. This implies that the zero counts follow a normal distribution and the sensor is stable.
Figure 14. Plot of maximum and rolling average zero counts over 21.5 hours
8.3. Further Improvements
It is possible to statistically investigate the area coverage of the delay sensor with regards to emitted laser power, scan duration, corresponding connections of the sensor circuitry (e.i, wires), and different tuning of the IDELAYs or any similar delay-based sensor against optical attacks. It would also be worthy to study the possible improvement of detection area coverage if multiple sensors are deployed in the target circuit.
9. Conclusion
In this paper, we presented the first fully digital and FPGA-compatible detection and response countermeasure, called LaserEscape, against optical probing attacks. LaserEscape leverages delay-based sensors to detect optical probing attempts and MTD strategies through the PR feature of conventional FPGAs and polymorphic circuits to respond to the attack. By repurposing the 1LUTsensor, deploying open-source bitstream manipulators, and designing LUT-based polymorphic gates, we built a real-time countermeasure to detect the laser beam and deceive the adversary. We demonstrated the resiliency of our approach by mounting optical probing attacks against target implementations on a 28 nm FPGA in two scenarios, namely key extraction and function extraction.
Acknowledgment
This effort was sponsored by NSF Grants CNS-2150123 and CNS-2150122.